# The emergence of galactic thin and thick discs across cosmic history

### v1

@@ 2.3 Point spread function
An accurate point spread function is essential for studying the intrinsic light distribution of galaxies, particularly for identifying the faint, thick disc component, superimposed by the brighter thin disc component (Comerón et al., 2018). We measured the effective point spread function (ePSF) from each mosaic image, following star selection methods (Faisst et al. 2022, see also Ito et al. 2024 and Zhuang & Shen 2024) [^2] We used the JWST PSF Pipeline and its forked version https://github.com/takafumi291/JPP.. For each field, we run sextractor to identify candidate stars with source properties class_star $>$ 0.8 and elongation $<$ 1.5. We select bright, unsaturated stars that lay on a horizontal locus in the magnitude-size plane, with a minimum signal-to-noise ratio (SNR) $>15$ and no brighter sources within 2". We then stacked the image cutouts of these selected stars using the Python package Photutils (Bradley et al., 2024; Anderson & King, 2000; Anderson, 2016), maintaining the identical pixel sampling of the images to include the exact pixelization effect.
The measured ePSF per mosaic captures the point source response influenced by telescope jittering (Morishita et al., 2024), the image drizzling process, and pixelization. This ePSF is expected to be broader than the simulated PSF by WebbPSF (Perrin et al., 2014), which represents the intrinsic PSF properties without these effects.
@@
where $\nu_{0}$ is the central luminosity density, and $h_{R}$ and $z_{0}$ represent the disc scale length and height, respectively. The $\operatorname{sech}^{2}$ profile, a solution for the self-gravitating isothermal sheet (Spitzer, 1942), has been widely used to effectively describe and approximate the vertical thickness of galactic discs (van der Kruit, 1988; Yoachim & Dalcanton, 2008). The generalized function, $\operatorname{sech}^{2/N}$, proposed by (van der Kruit & Searle, 1982) describes non-isothermal disc profiles. The model varies from the standard $\operatorname{sech}^{2}$ profile only near the midplane (van der Kruit & Searle, 1982), becoming either peaked or smoothed as $N$ approaches 1 or $\infty$, respectively. Despite these differences, all variants of the model asymptotically converge to the exponential profile $\exp(-z/z_{0})$ at large distance, $z$, from the midplane. In the limiting case where $N\rightarrow\infty$, the model matches the exponential profile $\exp(-z/z_{0})$ at all radii.
@@
For this study, we use the $\operatorname{sech}^{2}$ profile for consistency with existing literature. Detailed modelling of the midplane profile is beyond our scope, largely due to dust absorption and the limited resolution of the JWST. We also assume an inclination of $90^{\circ}$. Our population model predicts that 64% of our galaxies deviate from $90^{\circ}$ by no more than $7^{\circ}$ (see Section 2.2). Such a slight deviation from an inclination of $90^{\circ}$ introduces bias in the structural parameter measurements, which is known to be small (de Grijs et al., 1997). However, the impact depends on the intrinsic axial ratio of the disc and image quality (e.g., signal to noise ratio and instrumental resolution relative to the source extent). In Appendix C, we estimate potential biases introduced by inclination deviations. The median biases from this assumption (for single disc, thin disc, and thick disc components) vary with disc properties: (i) for disc scale radii, these are underestimated by 0.1%, 12%, 1% respectively; (ii) for disc scale-heights, there are overestimated by 17%, 27%, 9% respectively.
@@
First, we fit a single 3D disc model to each cutout JWST image for the galaxy sample at the selected waveband for maximal overlap with rest-frame $K_{s}$ band (e.g., F277W; $z<0.46$, F356W; $0.820.46<z<$ 0.82, F [^4] W; $0.82<z$) using the imfit package (Erwin, 2015). IMFIT projects the 3D luminosity along the line of sight onto 2D surface brightness in the sky, convolves with the user-supplied PSF, and then minimizes $\chi^{2}$ with the Levenberg–Marquardt algorithm. We set the lower boundary for the disc scale height to 0.2 pixels, based on our recovery experiment (see Appendix C.1). If the parameter reaches this boundary, we interpret it as a upper limit on the scale height. The best-fit single-disc model captures most of the galaxy’s total flux in the data. However, some galaxies show systematic residual patterns in a model-subtracted data map. The most notable are disc-like excess light off the midplane and compact bulge components in the centre of galaxies.
@@
Secondly, to account for these structures and measure their structural properties, we consider two additional components: 1) a Sérsic profile to account for the centrally concentrated light and 2) a second thicker 3D disc component to account for the disc-like excess light. For the first component, we allow the Sérsic index to range from 0.5 to 5, accommodating the range from classical bulges to disc-like bulges (Kormendy & Kennicutt, 2004).
@@
To find the best structural fit for each galaxy, we fit a series of increasingly complex models that include the second disc and central Sérsic component. These include (i) the single disc model described above; (ii) two disc model (thin+thick); (iii) a disc+bulge model; and (iv) a two disc (thin+thick) + bulge model. To avoid local minima in the $\chi^{2}$ landscape for the multiple component fits, we repeated Levenberg-Marquard optimization 15 times with initial starting points randomly drawn from a conservative wide range of parameters based on the single disc fits. To avoid erroneous solutions for the bulge being fit to the disc structure, and vice versa for the model including the bulge component, the $\chi^{2}$ minimum solution was bounded by a constraint: the bulge effective radius is smaller than both 2 kpc and the disc scale length.
@@
To ensure robust fits with well-behaved components motivated by data, we use the following procedures. We determine the necessity for additional components for each galaxy by comparing the Bayesian Information Criterion (BIC) across models. We assigned the galaxies into 4 models, from the simplest to more complex models as described above. A more complex model is justified if the BIC improvement exceeds 15. We visually inspect the fit results using the model-subtracted residuals, checking each structural component fits the intended structure.
@@
Although the BIC classification is generally robust, we reclassify 6 galaxies from ‘two discs + bulge’ to ‘a disc + bulge’, and 1 from ‘two discs’ to ‘single disc’, based on the visual insignificance of the thick disc component [^3] The reclassified categories are the second best model of these galaxies with the smallest BIC differences among other galaxies in the pre-reclassified category.. One galaxy initially categorized as "single + bulge" was adjusted to ‘single disc’, while three others were reclassified to ‘two discs’ (2) and ‘two discs + bulge’ (1) because the bulge component fits the visually significant disc feature. These 11 re-classifications are only a minor subset of the 111 galaxies -- eliminating them instead did not alter our conclusions of the paper.
@@
Figure 4: An example fit of a galaxy (ID=1) in our sample, with ‘two discs + bulge’ model being the best model according to our criteria (Section 2.4). The three groups of panels from top to bottom display the best fit with (a) a single disc, (b) two discs, and (c) two discs + a bulge, respectively. Each group of panels show from left to right the data, model, and residual in the top row and the vertical surface brightness profiles at different radii in the bottom. Each profile is extracted at the radius indicated above it. Assuming the two-fold symmetry of the disc galaxies, we show the average profile of four quadrants. The galaxy’s redshift $z$, $\chi^{2}_{\nu}$ of each fit and derived disc parameters are also shown. Notably, there remains significant excess light above the disc midplane for the single disc model, which is eliminated by adding a second disc component. The central excess is accounted for by two discs + bulge models, but even ignoring it entirely, the derived parameters do not change by more than 20%, consistent with the conservative systematic error of our sample (Appendix C).
Figure 4 shows an example of a galaxy (ID=1) identified as ‘two discs + bulge’ with fits for a single disc (top), two discs (middle), and two discs + a bulge (bottom) models. For this galaxy, the ‘single disc + bulge’ model provides a similar fit to two disc models as the bulge becomes the thin disc component. So, we omit it to avoid redundancy.
@@
As the model is not always perfect, there are higher-order structures unmodelled, such as the outer truncation of the disc, and disc substructure (bar end or clump, see Fig. 4 for residual at $-10$ kpc). Although not recognizable in our chosen bands, dust extinction may bias our results. Such challenges are noted in previous works for nearby galaxies (Yoachim & Dalcanton, 2006). To test the robustness of the results, we have run additional tests described in Appendix C.3. This includes using masks of different regions of the galaxies (e.g., disc midplane, centre, outside typical truncation radius) and systematics (e.g., inclination, potential dust extinction). In short, our conservative estimate of systematic and statistical uncertainty is about 15% for the scale radii and roughly 30% for scale heights. With experiments where the bulge region is masked rather than modelled, we show the bulge structure does not influence the measured properties of the discs.
@@ 3.1.1 Existence of radial/vertical size - mass correlation
In Figs 5 and 6, we show the power-law fits for each population, where the length and height are proportional to the power of stellar masses, $M_{*}^{\beta}$. Table 1 summarizes the measured slope $\beta$ and intercept $\alpha$ with the Spearman’s rank correlation coefficients and associated p-values against the null hypothesis that there is no correlation between the parameters. To find the best fits, robust to outliers, we used LTSLINEFIT (Cappellari et al., 2013), the implementation of the least trimmed squares (LTS) approach (Rousseeuw & Leroy, 1987; Rousseeuw & Driessen, 2006). The method effectively finds the global minima of the $\chi^{2}$ computed for all possible subsets of the data excluding potential outliers up to half of the sample. This method can also take into account the uncertainty for all coordinates. We used the statistical uncertainty for our disc length and height measurements and 0.2 dex uncertainties for stellar masses (Schreiber et al., 2011).
@@ 3.2.1 Single discs
The increasing trend in $z_{0}$ toward the present is consistent with similar studies using local galaxies (Yoachim & Dalcanton, 2006), but contrasts with the declining trend reported by Lian et al. (2024). This discrepancy can be attributed to the sequential formation of the thick disc followed by the thin disc and the use of shorter wavebands in Lian et al. (2024), which are sensitive to younger stellar populations. Consequently, their measurements capture the thickness associated with the ongoing formation of thick discs at earlier times and the subsequent formation of thin discs at later times. This is further clarified in the next subsection (Sec. 3.2.2), which examines the evolution of thin and thick discs.
No median evolution is seen for the ratio of scale length to scale height ($h_{R}/z_{0}$, Fig. 7 right). This suggests that the discs at all of our explored epochs, $\tau_{\mathrm{lookback}}\sim 1.6-11.4$ Gyrs, have already developed geometrically similar structures to present-day galaxies (as denoted by the grey band at $z=0$). Mechanisms responsible for radial and vertical height growth maintain the geometrical proportion of the discs.
@@ 3.2.2 Thick and thin discs
Fig. 8 shows equivalent plots to those presented in Fig. 7, i.e. disc properties as a function of lookback time, but for the double disc subsample, showing both thin and thick discs. The measured scale heights and scale lengths for discs are almost constant as a function of lookback time after correcting the $M_{*}$ dependency. This indicates that, similar to single discs, there is no significant structural evolution in thin and thick discs.
One notable exception is a mild decrease in scale lengths for the thin disc components towards late times. This trend is of interest as thin discs may be preferentially affected by scattering processes that reduce their scale radius by decreasing angular momentum. However, measuring the scale length of thinner discs is subject to larger uncertainties (see Appendix C), and low number statistics at the bin at the latest lookback time.
@@
The median evolution of disc structural parameters discussed in this section reflects evolution in the overall population rather than the specific evolutionary tracks of individual galaxies. As galaxies and discs grow in mass, both their radii ($h_{R}$) and heights ($z_{0}$) tend to increase on avarage, aligning with the scaling relations in Figs. 5 and 6. An increase of about two orders of magnitude in mass can roughly triple these sizes. In contrast, the redshift evolution affects the sizes by a factor of at most 1.5, contributing to the observed scatter in the scaling relation. The relatively constant or mildly evolving mass-corrected medians shown in Figs. 7 and 8 indicate that the fundamental scaling relations remain largely unchanged over time.
3.3 Emerging thin discs at later epochs of universe

### v2

@@ 2.3 Point spread function
An accurate PSF is essential for studying the intrinsic light distribution of galaxies, particularly for identifying the faint, thick disc component, superimposed by the brighter thin disc component (Comerón et al., 2018). We measured the effective PSF (ePSF) from each mosaic image, following star selection methods (Faisst et al. 2022, see also Ito et al. 2024 and Zhuang & Shen 2024) [^2] We used the JWST PSF Pipeline and its forked version https://github.com/takafumi291/JPP.. For each field, we run SExtractor to identify candidate stars with source properties class_star $>$ 0.8 and elongation $<$ 1.5. We select bright, unsaturated stars that lay on a horizontal locus in the magnitude-size plane, with a minimum signal-to-noise ratio (SNR) >20absent20>20> 20333This corresponds to AB magnitudes brighter than 22.5-23.5 depending on the filter and field. and no brighter sources within 2". We then stacked the image cutouts of these selected stars using the Python package Photutils (Bradley et al., 2024; Anderson & King, 2000; Anderson, 2016), maintaining the identical pixel sampling of the images to include the exact pixelization effect.
The measured ePSF per mosaic captures the point source response influenced by telescope jittering (Morishita et al., 2024), the image drizzling process, and pixelization. This ePSF is expected to be broader than the simulated PSF by WebbPSF (Perrin et al., 2014), which represents the intrinsic PSF properties without these effects. By fitting a Gaussian function to the ePSF, we measured the ePSF FWHM of 0.131", 0.146", and 0.166" for F277W, F356W, and F [^4] W, respectively. The measured values do not change significantly across observations. Because the ePSF is only roughly approximated by a Gaussian, the measured ePSF widths depend on the pixel sampling; here, we used a pixel sampling size of 0.40". In physical scales, these FWHM corresponds to 0.3 kpc - 1.4 kpc with a median of 1.1 kpc in our galaxy sample, given the galaxy distance and band used. The limitation on the smallest measurable structural scale from the finite PSF width is evaluated in Sec. 2.4 and in Appendix C.
@@
where $\nu_{0}$ is the central luminosity density, and $h_{R}$ and $z_{0}$ represent the disc scale length and height, respectively. The $\operatorname{sech}^{2}$ profile, a solution for the self-gravitating isothermal sheet (Spitzer, 1942), has been widely used to effectively describe and approximate the vertical thickness of galactic discs (van der Kruit, 1988; Yoachim & Dalcanton, 2008). The generalized function, $\operatorname{sech}^{2/N}$, proposed by van der Kruit & Searle (1982) describes non-isothermal disc profiles. The model varies from the standard $\operatorname{sech}^{2}$ profile only near the midplane (van der Kruit & Searle, 1982), becoming either peaked or smoothed as $N$ approaches 1 or $\infty$, respectively. Despite these differences, all variants of the model asymptotically converge to the exponential profile $\exp(-z/z_{0})$ at large distance, $z$, from the midplane. In the limiting case where $N\rightarrow\infty$, the model matches the exponential profile $\exp(-z/z_{0})$ at all radii.
@@
For this study, we use the $\operatorname{sech}^{2}$ profile for consistency with existing literature. Detailed modelling of the midplane profile is beyond our scope, largely due to dust absorption and the limited resolution of JWST. We also assume an inclination of $90^{\circ}$. Our population model predicts that 64% of our galaxies deviate from $90^{\circ}$ by no more than $7^{\circ}$ (see Section 2.2). Such a slight deviation introduces bias in the structural parameter measurements, which is known to be small (de Grijs et al., 1997). However, the impact depends on the intrinsic axial ratio of the disc and image quality (e.g., signal to noise ratio and instrumental resolution relative to the source extent). Section 2.5 discusses systematic uncertainties from inclination and other factors for structural measurements.
To model galaxy surface brightness distributions and measure the structural properties, we use the imfit package (Erwin, 2015) throughout the paper. imfit projects the 3D luminosity along the line of sight onto 2D surface brightness in the sky, convolves with the user-supplied PSF, and then minimizes $\chi^{2}$ with the Levenberg–Marquardt algorithm. It takes the image, variance, and mask (to define the fitting region) as inputs and outputs the best-fit parameters and fit metrics such as the best fit $\chi^{2}$ and Bayesian Information Criterion (BIC). Additionally, imfit allows for the superposition of multiple components such as a second disc or 2D Sérsic profile, which we employed in later stages.
@@
First, we fit a single 3D disc model to each cutout JWST image for the galaxy sample at the selected waveband for maximal overlap with rest-frame $K_{s}$ band (e.g., F277W; $z<0.46$, F356W; $0.820.46<z<$ 0.82, F [^4] W; $0.82<z$). We set the lower boundary for the disc scale height to 0.2 pixels, based on our recovery experiment (see Appendix C.1). If the parameter reaches this boundary, we interpret it as a upper limit on the scale height. The best-fit single-disc model captures most of the galaxy’s total flux in the data. However, some galaxies show systematic residual patterns in a model-subtracted data map. The most notable are disc-like excess light off the midplane and compact bulge components in the centre of galaxies.
@@
Secondly, to account for these structures and measure their structural properties, we consider two additional components: 1) a 2D Sérsic profile to account for the centrally concentrated light and 2) a second thicker 3D disc component to account for the disc-like excess light. For the first component, we allow the Sérsic index to range from 0.5 to 5, accommodating the range from classical bulges to disc-like bulges (Kormendy & Kennicutt, 2004).
@@
To find the best structural fit for each galaxy, we fit a series of increasingly complex models that include the second disc and central Sérsic component using imfit. These include (i) the single disc model described above; (ii) two disc model (thin+thick); (iii) a disc+bulge model; and (iv) a two disc (thin+thick) + bulge model. To avoid local minima in the $\chi^{2}$ landscape for the multiple component fits, we repeated Levenberg-Marquard optimization 15 times with initial starting points randomly drawn from a conservative wide range of parameters based on the single disc fits. To avoid erroneous solutions for the bulge being fit to the disc structure, and vice versa for the model including the bulge component, the $\chi^{2}$ minimum solution was bounded by a constraint: the bulge effective radius is smaller than both 2 kpc and the disc scale length.
@@
To ensure robust fits with well-behaved components motivated by data, we use the following procedures. We determine the necessity for additional components for each galaxy by comparing the BIC across models. We assigned the galaxies into 4 models, from the simplest to more complex models as described above. A more complex model is justified if the BIC improvement exceeds 15. We visually inspect the fit results using the model-subtracted residuals, checking each structural component fits the intended structure.
@@
Although the BIC classification is generally robust, we reclassify 6 galaxies from ‘two discs + bulge’ to ‘a disc + bulge’, and 1 from ‘two discs’ to ‘single disc’, based on the visual insignificance of the thick disc component. [^4] In all 6 cases where classification changed based on visual inspection, the adopted classification corresponds to the second-best model when evaluating the BIC. Additionally, these 6 galaxies had some of the smallest differences in BIC between the best and second-best models among the full sample of galaxies. One galaxy initially categorized as "single + bulge" was adjusted to ‘single disc’, while three others were reclassified to ‘two discs’ (2) and ‘two discs + bulge’ (1) because the bulge component fits the visually significant disc feature. These 11 re-classifications are only a minor subset of the 111 galaxies -- eliminating them instead did not alter our conclusions of the paper.
@@
Figure 4: An example fit of a galaxy (ID=1) in our sample, with ‘two discs + bulge’ model being the best model according to our criteria (Section 2.4). The three groups of panels from top to bottom display the best fit with (a) a single disc, (b) two discs, and (c) two discs + a bulge, respectively. Each group of panels show from left to right the data, model, and residual in the top row and the vertical surface brightness profiles at different radii in the bottom. All images and profiles are PSF-convolved. Each profile is extracted at the radius indicated above it. Assuming the two-fold symmetry of the disc galaxies, we show the average profile of four quadrants. The galaxy’s redshift $z$, $\chi^{2}_{\nu}$ of each fit and derived disc parameters are also shown. Notably, there remains significant excess light above the disc midplane for the single disc model, which is eliminated by adding a second disc component. The central excess is accounted for by two discs + bulge models, but even ignoring it entirely, the derived parameters do not change by more than 20%, consistent with the conservative systematic error of our sample (Appendix C).
Figure 4 shows an example galaxy (ID=1) best fitted with a ‘two discs + bulge’ model alongside fits with a single disc (top), two discs (middle), and two discs + a bulge (bottom) models. For this galaxy, the ‘single disc + bulge’ model provides a similar fit to two disc models as the bulge becomes the thin disc component. We omit it here to avoid redundancy. Fitting a single disc model leaves systematic flux excesses above the disc midplane, clearly seen in both the residual image and vertical profile (Fig. 4a). Adding a second disc component successfully accounts for this excess light (Fig. 4b). The convolved profiles reveal diversity amongst the sample. Some galaxies are dominated by either thin or thick discs, while others transition between thin/thick dominance at a certain height. However, it is difficult to assess the relative dominance of the thin and thick discs in the PSF-convolved vertical profiles shown in Fig. 4. PSF wings can cause the thin disc light to extend to large heights. Characterising the PSF-deconvolved vertical profile of each component will be left for future work, which will require verifying the deconvolved profiles and assessing the uncertainties propagated from both the assumed model and statistical noise.
2.5 Uncertainties for structural measurements
JWST data allows us to measure the disc structural properties with high precision. We estimated statistical uncertainties using Levenberg-Marquardt optimization in IMFIT and verified them by bootstrap resampling, refitting the resampled data 500 times. We present our results with these statistical uncertainties throughout the paper. We also separately estimate systematic uncertainty arising from our assumptions and simplifications, as detailed in Appendix C, to help readers interpret our measurements and their limitations.
@@ 2.5 Uncertainties for structural measurements
As the model is not always perfect, higher-order structures remain unmodelled, such as the outer truncation of the disc, and disc substructure (bar end or clump, see Fig. 4 for residual at $-10$ kpc). Although dust extinction is not evident in our chosen bands, it may also bias our results. Such challenges are noted in previous works for nearby galaxies (Yoachim & Dalcanton, 2006). To test the robustness of our results, we have run additional tests that includes masking different regions of the galaxies (e.g., disc midplane, centre, outside typical truncation radius) and systematic effect (e.g., inclination, potential dust extinction). In summary, systematic uncertainty, mainly arising from deviations in the inclination from 90, results in median underestimation of 0.1%, 12%, 1% for scale radii and median overestimations of 17%, 27%, 9% for the scale heights for single, thin, thick disc respectively. Experiments masking the bulge region rather than modelling it confirm that the bulge structure does not influence the measured properties of the discs. The systematic effects are not significant compared to the reported trends in Section 3 and as such do not affect our conclusions.
@@ 3.1.1 Existence of radial/vertical size - mass correlation
In Figs 5 and 6, we show power-law fits for each population, where the length and height are proportional to the power of stellar masses, $M_{*}^{\beta}$. Table 1 summarizes the measured slope $\beta$ and intercept $\alpha$ with the Spearman’s rank correlation coefficients and associated p-values against the null hypothesis that there is no correlation between the parameters. To find the best fits, robust to outliers, we used LTSLINEFIT (Cappellari et al., 2013), the implementation of the least trimmed squares (LTS) approach (Rousseeuw & Leroy, 1987; Rousseeuw & Driessen, 2006). The method effectively finds the global minima of the $\chi^{2}$ computed for all possible subsets of the data excluding potential outliers up to half of the sample. This method can also take into account the uncertainty for all coordinates. We used the statistical uncertainty for our disc length and height measurements and 0.2 dex uncertainties for stellar masses (Schreiber et al., 2011).
@@ 3.2.1 Single discs
The increasing trend in $z_{0}$ toward the present is consistent with similar studies using local galaxies (Yoachim & Dalcanton, 2006), but contrasts with the declining trend reported by Lian & Luo (2024). This discrepancy can be attributed to the sequential formation of the thick disc followed by the thin disc and the use of shorter wavebands in Lian & Luo (2024), which are sensitive to younger stellar populations. Consequently, their measurements capture the thickness associated with the ongoing formation of thick discs at earlier times and the subsequent formation of thin discs at later times. This is further clarified in the next subsection (Sec. 3.2.2), which examines the evolution of thin and thick discs.
No median evolution is seen for the ratio of scale length to scale height ($h_{R}/z_{0}$, Fig. 7 right). This suggests that the discs at all of our explored epochs, $\tau_{\mathrm{lookback}}\sim 1.6-11.4$ Gyrs, have already developed geometrically similar structures to present-day galaxies (as denoted by the grey band at $z=0$).
@@ 3.2.2 Thick and thin discs
Fig. 8 shows equivalent plots to those presented in Fig. 7, i.e. disc properties as a function of lookback time, but for the double disc subsample, showing both thin and thick discs. The measured scale lengths and heights for discs are almost constant as a function of lookback time after correcting the $M_{*}$ dependency. This indicates that there is no significant structural evolution in thin and thick discs.
One notable exception is a mild decrease in scale lengths for the thin disc components towards late times. This trend is of interest as thin discs may be preferentially affected by scattering processes that reduce their scale radius by decreasing angular momentum, or because we observe the effect of reduced pressure support in the thin disc, causing it to shrink (see more discussion in Sec. 4.3). However, measuring the scale length of thinner discs is subject to larger uncertainties (see Appendix C), and low number statistics at the bin at the latest lookback time.
@@
3.2.3 Interpreting evolutionary trends in disc formation
Single discs show a mild increase in both scale length ($h_{R}$) and height ($z_{0}$) toward the present, whereas the thin and thick discs in two-disc galaxies show no significant evolution. It is important to note that the median evolution of disc structural parameters at a given galaxy mass does not reflect the specific evolutionary tracks of individual galaxies – galaxies in different time bins are not the same progenitor populations. With the mass dependency already removed, the evolution in radius and height is primarily determined by angular momentum and vertical energy, respectively.
In the context of cosmic downsizing, for galaxies of the same mass, those observed at early times formed rapidly (Behroozi et al., 2013) and completed their thick disc formation earlier (Comerón, 2021) compared to galaxies observed at later times. The mild increase in $h_{R}$ at given galaxy mass can be attributed to halo evolution in the $\Lambda$CDM universe, where halo virial radii grow larger toward the present. Gas accreting from more extended regions carries larger specific angular momentum, so discs that form later tend to be larger (Mo et al., 1998). Meanwhile, lower halo concentrations at earlier epochs (Bullock et al., 2001) partially offsets this effect, resulting in a modest overall size increase (Somerville et al., 2008). Additional crucial processes, such as outflows removing low–angular momentum material (Brook et al., 2011; Guedes et al., 2011), radial migration that move stars to outer radii on average (Minchev et al., 2012), and mergers also affect disc sizes (Governato et al., 2009). The increase in disc thickness $z_{0}$ is expected because stars do not cool and are continuously thickened by satellite perturbations; consequently, discs observed at later times have more chance to be heated by such events. Moreover, for a given velocity dispersion, a disc with a larger scale length at the same mass has a lower surface mass density, leading to a larger scale height in gravitational equilibrium. The absence of an increase in $z_{0}$ for the thick disc in two-disc galaxies (middle panel of Fig. 8) could be due to the presence of the thin disc. Simulations have shown that gas in a thin disc can reduce heating from minor mergers and thin disc growth can lead to the adiabatic vertical contraction of the thick disc (Moster et al., 2010). In our disc sample, the mechanisms responsible for radial and vertical height growth seem to maintain the geometrical proportion of the discs (see constant $h_{R}/z_{0}$ in Figs. 7 and 8).
For evolutionary tracks of individual galaxies, it is more informative to examine their size and mass relations. As galaxies and discs grow in mass, both their radii ($h_{R}$) and heights ($z_{0}$) tend to increase on average, aligning with the scaling relations in Figs. 5 and 6. An increase of about two orders of magnitude in mass can roughly triple these sizes. In contrast, the redshift evolution affects the sizes by a factor of at most 1.5, contributing to the observed scatter in the scaling relation. The relatively constant or mildly evolving mass-corrected medians shown in Figs. 7 and 8 indicate that the fundamental scaling relations remain largely unchanged over time.
3.3 Emerging thin discs at later epochs